in two regular polygons having the same number of sides find the area of smaller if sides have length 4 and 12 and area of larger is 10 260
1
Expert's answer
2020-05-13T18:38:11-0400

The area of ​​a regular polygon with the n number of sides n:

S= ((n*a2 /4)*сot(π/n)

S1=?

S2=10260

а1=4

а2=12

S1/S2 = a12/a22

S1/10260=16/144

S1=10260*16/144=1140

Answer S1=1140